A cylindrical can of radius 5 cm and height 8 cm is used to pour water into a larger cylinder of radius 20 cm and height 2m. How many times must this be done to fill the larger cylinder? plz help me. Thanks in advance.
Answer by jim_thompson5910(35256)

Volume of smaller cylinder:

V = pi*r^2*h

V = pi*5^2*8

V = 200pi

The volume of the small cylinder is 200pi cubic cm
-------------------------------------------------------

Volume of larger cylinder:

V = pi*r^2*h

V = pi*20^2*2

V = 800pi

The volume of the larger cylinder is 800pi cubic cm

-------------------------------------------------------

Divide the two (large/small)

(800pi/200pi) = 800/200 = 4

So you need to pour all the contents of the small cylinder into the larger one 4 times.

Let me know if you need more help or if you need me to explain a step in more detail.
Feel free to email me at jim_thompson5910@hotmail.com
or you can visit my website here: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim